In triangle ABC, b=11, c=18, and angle B=40 degrees. solve the triangle.
Answer by venugopalramana(3286)

In triangle ABC, b=11, c=18, and angle B=40 degrees. solve the triangle.
b/SIN B =c/SIN C..HENCE
SIN C = c*SIN B / b
SIN C = 18*SIN 40 /11= 1.051390552
THE PROBLEM IS NOT CORRECT .PLEASE CHECK UP THE VALUES OF b AND c